# MIRACLE II: Unveiling the multi-phase gas interplay in the circumnuclear region of NGC 1365 via multi-cloud modeling

### v3

@@ Acknowledgements.
- de Vaucouleurs et al. (1991) de Vaucouleurs, G., de Vaucouleurs, A., Corwin, Jr., H. G., et al. 1991, Third Reference Catalogue of Bright Galaxies
- Decleir et al. (2022) Decleir, M., Gordon, K. D., Andrews, J. E., et al. 2022, ApJ, 930, 15
- Dopita et al. (2006) Dopita, M. A., Fischera, J., Sutherland, R. S., et al. 2006, ApJS, 167, 177
- Dudik et al. (2007) Dudik, R. P., Weingartner, J. C., Satyapal, S., et al. 2007, ApJ, 664, 71
- Fabian (2012) Fabian, A. C. 2012, ARA&A, 50, 455
- Feltre et al. (2023) Feltre, A., Gruppioni, C., Marchetti, L., et al. 2023, A&A, 675, A74
- Ferland et al. (1998) Ferland, G. J., Korista, K. T., Verner, D. A., et al. 1998, PASP, 110, 761
- Feruglio et al. (2010) Feruglio, C., Maiolino, R., Piconcelli, E., et al. 2010, A&A, 518, L155
- Fiore et al. (2017) Fiore, F., Feruglio, C., Shankar, F., et al. 2017, A&A, 601, A143
- Fitzpatrick et al. (2019) Fitzpatrick, E. L., Massa, D., Gordon, K. D., Bohlin, R., & Clayton, G. C. 2019, ApJ, 886, 108
- Fluetsch et al. (2021) Fluetsch, A., Maiolino, R., Carniani, S., et al. 2021, MNRAS, 505, 5753
- Fluetsch et al. (2019) Fluetsch, A., Maiolino, R., Carniani, S., et al. 2019, MNRAS, 483, 4586
- Forbes & Norris (1998) Forbes, D. A. & Norris, R. P. 1998, MNRAS, 300, 757
- Förster Schreiber et al. (2019) Förster Schreiber, N. M., Übler, H., Davies, R. L., et al. 2019, ApJ, 875, 21
- Gao et al. (2021) Gao, Y., Egusa, F., Liu, G., et al. 2021, ApJ, 913, 139
- García-Bernete et al. (2024) García-Bernete, I., Alonso-Herrero, A., Rigopoulou, D., et al. 2024, A&A, 681, L7
- García-Burillo et al. (2019) García-Burillo, S., Combes, F., Ramos Almeida, C., et al. 2019, A&A, 632, A61
- Garofali et al. (2024) Garofali, K., Basu-Zych, A. R., Johnson, B. D., et al. 2024, ApJ, 960, 13
- Gordon (2024) Gordon, K. 2024, The Journal of Open Source Software, 9, 7023
- Gordon et al. (2009) Gordon, K. D., Cartledge, S., & Clayton, G. C. 2009, ApJ, 705, 1320
- Gordon et al. (2023) Gordon, K. D., Clayton, G. C., Decleir, M., et al. 2023, ApJ, 950, 86
- Gordon et al. (2021) Gordon, K. D., Misselt, K. A., Bouwman, J., et al. 2021, ApJ, 916, 33
- Groves et al. (2004) Groves, B. A., Dopita, M. A., & Sutherland, R. S. 2004, ApJS, 153, 75
- Gutkin et al. (2016) Gutkin, J., Charlot, S., & Bruzual, G. 2016, MNRAS, 462, 1757
- Hao et al. (2009) Hao, L., Wu, Y., Charmandaris, V., et al. 2009, ApJ, 704, 1159
- Harrison et al. (2016) Harrison, C. M., Alexander, D. M., Mullaney, J. R., et al. 2016, MNRAS, 456, 1195
- Harrison et al. (2014) Harrison, C. M., Alexander, D. M., Mullaney, J. R., & Swinbank, A. M. 2014, MNRAS, 441, 3306
- Harrison & Ramos Almeida (2024) Harrison, C. M. & Ramos Almeida, C. 2024, Galaxies, 12, 17
- Hatziminaoglou et al. (2015) Hatziminaoglou, E., Zwaan, M., Andreani, P., et al. 2015, The Messenger, 162, 24
- Hermosa Muñoz et al. (2024a) Hermosa Muñoz, L., Alonso-Herrero, A., Labiano, A., et al. 2024a, arXiv e-prints, arXiv:2412.14707
- Hermosa Muñoz et al. (2024b) Hermosa Muñoz, L., Alonso-Herrero, A., Pereira-Santaella, M., et al. 2024b, A&A, 690, A350
- Hill & Zakamska (2014) Hill, M. J. & Zakamska, N. L. 2014, MNRAS, 439, 2701
- Inami et al. (2013) Inami, H., Armus, L., Charmandaris, V., et al. 2013, ApJ, 777, 156
- Jacobs et al. (2009) Jacobs, B. A., Rizzi, L., Tully, R. B., et al. 2009, AJ, 138, 332
- Jones & Jones (1980) Jones, J. E. & Jones, B. J. T. 1980, MNRAS, 191, 685
- Kakkad et al. (2018) Kakkad, D., Groves, B., Dopita, M., et al. 2018, A&A, 618, A6
- Kakkad et al. (2020) Kakkad, D., Mainieri, V., Vietri, G., et al. 2020, A&A, 642, A147
- Kewley et al. (2001) Kewley, L. J., Dopita, M. A., Sutherland, R. S., Heisler, C. A., & Trevena, J. 2001, ApJ, 556, 121
- Kewley et al. (2006) Kewley, L. J., Groves, B., Kauffmann, G., & Heckman, T. 2006, MNRAS, 372, 961
- Kewley et al. (2013) Kewley, L. J., Maier, C., Yabe, K., et al. 2013, ApJ, 774, L10
- Kristen et al. (1997) Kristen, H., Jorsater, S., Lindblad, P. O., & Boksenberg, A. 1997, A&A, 328, 483
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Labiano et al. (2021) Labiano, A., Argyriou, I., Álvarez-Márquez, J., et al. 2021, A&A, 656, A57
- Laor (1998) Laor, A. 1998, ApJ, 496, L71
- Law et al. (2023) Law, D. R., E. Morrison, J., Argyriou, I., et al. 2023, AJ, 166, 45
- Lee et al. (2023) Lee, J. C., Sandstrom, K. M., Leroy, A. K., et al. 2023, ApJ, 944, L17
- Lee et al. (2022) Lee, J. C., Whitmore, B. C., Thilker, D. A., et al. 2022, ApJS, 258, 10
- Lindblad et al. (1996) Lindblad, P. O., Hjelm, M., Hoegbom, J., et al. 1996, A&AS, 120, 403
- Liu et al. (2023) Liu, D., Schinnerer, E., Cao, Y., et al. 2023, ApJ, 944, L19
- Luridiana et al. (2015) Luridiana, V., Morisset, C., & Shaw, R. A. 2015, A&A, 573, A42
- Lutz et al. (2020) Lutz, D., Sturm, E., Janssen, A., et al. 2020, A&A, 633, A134
- Marasco et al. (2020) Marasco, A., Cresci, G., Nardini, E., et al. 2020, A&A, 644, A15
- Marconcini et al. (2025a) Marconcini, C., Feltre, A., Lamperti, I., et al. 2025a, A&A, 701, A113
- Marconcini et al. (2025b) Marconcini, C., Marconi, A., Cresci, G., et al. 2025b, Nature Astronomy, 9, 907
- Marconcini et al. (2023) Marconcini, C., Marconi, A., Cresci, G., et al. 2023, A&A, 677, A58
- Marconi et al. (2024) Marconi, A., Amiri, A., Feltre, A., et al. 2024, A&A, 689, A78
- Martínez-Paredes et al. (2023) Martínez-Paredes, M., Bruzual, G., Morisset, C., et al. 2023, MNRAS, 525, 2916
- Mingozzi et al. (2019) Mingozzi, M., Cresci, G., Venturi, G., et al. 2019, A&A, 622, A146
- Mingozzi et al. (2025) Mingozzi, M., Garcia del Valle-Espinosa, M., James, B. L., et al. 2025, ApJ, 985, 253
- Morrison et al. (2023) Morrison, J. E., Dicken, D., Argyriou, I., et al. 2023, PASP, 135, 075004
- Mulcahey et al. (2022) Mulcahey, C. R., Leslie, S. K., Jackson, T. M., et al. 2022, A&A, 665, A144
- Nicholls et al. (2017) Nicholls, D. C., Sutherland, R. S., Dopita, M. A., Kewley, L. J., & Groves, B. A. 2017, MNRAS, 466, 4403
- Osterbrock & Ferland (2006) Osterbrock, D. E. & Ferland, G. J. 2006, Astrophysics of gaseous nebulae and active galactic nuclei
- Patapis et al. (2024) Patapis, P., Argyriou, I., Law, D. R., et al. 2024, A&A, 682, A53
- Perna et al. (2024) Perna, M., Arribas, S., Lamperti, I., et al. 2024, A&A, 690, A171
- Perrin et al. (2014) Perrin, M. D., Sivaramakrishnan, A., Lajoie, C.-P., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9143, Space Telescopes and Instrumentation 2014: Optical, Infrared, and Millimeter Wave, ed. J. M. Oschmann, Jr., M. Clampin, G. G. Fazio, & H. A. MacEwen, 91433X
- Ramos Almeida et al. (2022) Ramos Almeida, C., Bischetti, M., García-Burillo, S., et al. 2022, A&A, 658, A155
- Ramos Almeida et al. (2025) Ramos Almeida, C., Garcia-Bernete, I., Pereira-Santaella, M., et al. 2025, arXiv e-prints, arXiv:2504.01595
- Richardson et al. (2022) Richardson, C. T., Simpson, C., Polimera, M. S., et al. 2022, ApJ, 927, 165
- Richings & Faucher-Giguère (2018a) Richings, A. J. & Faucher-Giguère, C.-A. 2018a, MNRAS, 478, 3100
- Richings & Faucher-Giguère (2018b) Richings, A. J. & Faucher-Giguère, C.-A. 2018b, MNRAS, 474, 3673
- Rieke et al. (2015) Rieke, G. H., Wright, G. S., Böker, T., et al. 2015, PASP, 127, 584
- Riffel et al. (2020) Riffel, R. A., Zakamska, N. L., & Riffel, R. 2020, MNRAS, 491, 1518
- Risaliti et al. (2005) Risaliti, G., Bianchi, S., Matt, G., et al. 2005, ApJ, 630, L129
- Sakamoto et al. (2007) Sakamoto, K., Ho, P. T. P., Mao, R.-Q., Matsushita, S., & Peck, A. B. 2007, ApJ, 654, 782
- Sanchez (2009) Sanchez, R. Z. 2009, PhD thesis, Rutgers University, New Jersey
- Sandqvist et al. (1995) Sandqvist, A., Joersaeter, S., & Lindblad, P. O. 1995, A&A, 295, 585
- Schinnerer et al. (2023) Schinnerer, E., Emsellem, E., Henshaw, J. D., et al. 2023, ApJ, 944, L15
- Sharp & Bland-Hawthorn (2010) Sharp, R. G. & Bland-Hawthorn, J. 2010, ApJ, 711, 818
- Speranza et al. (2024) Speranza, G., Ramos Almeida, C., Acosta-Pulido, J. A., et al. 2024, A&A, 681, A63
- Stanway & Eldridge (2018) Stanway, E. R. & Eldridge, J. J. 2018, MNRAS, 479, 75
- Stevens et al. (1999) Stevens, I. R., Forbes, D. A., & Norris, R. P. 1999, MNRAS, 306, 479
- Teuben et al. (1986) Teuben, P. J., Sanders, R. H., Atherton, P. D., & van Albada, G. D. 1986, MNRAS, 221, 1
- Tozzi et al. (2021) Tozzi, G., Cresci, G., Marasco, A., et al. 2021, A&A, 648, A99
- Ulivi et al. (2025) Ulivi, L., Perna, M., Lamperti, I., et al. 2025, A&A, 693, A36
- Vazdekis et al. (2016) Vazdekis, A., Koleva, M., Ricciardelli, E., Röck, B., & Falcón-Barroso, J. 2016, MNRAS, 463, 3409
- Veilleux et al. (2005) Veilleux, S., Cecil, G., & Bland-Hawthorn, J. 2005, ARA&A, 43, 769
- Veilleux & Osterbrock (1987) Veilleux, S. & Osterbrock, D. E. 1987, ApJS, 63, 295
- Venturi et al. (2021) Venturi, G., Cresci, G., Marconi, A., et al. 2021, A&A, 648, A17
- Venturi et al. (2017) Venturi, G., Marconi, A., Mingozzi, M., et al. 2017, Frontiers in Astronomy and Space Sciences, 4, 46
- Venturi et al. (2018) Venturi, G., Nardini, E., Marconi, A., et al. 2018, A&A, 619, A74
- Venturi et al. (2023) Venturi, G., Treister, E., Finlez, C., et al. 2023, A&A, 678, A127
- Véron-Cetty & Véron (2006) Véron-Cetty, M. P. & Véron, P. 2006, A&A, 455, 773
- Weaver et al. (2010) Weaver, K. A., Meléndez, M., Mushotzky, R. F., et al. 2010, ApJ, 716, 1151
- Woo et al. (2016) Woo, J.-H., Bae, H.-J., Son, D., & Karouzos, M. 2016, ApJ, 817, 108
- Wootten & Thompson (2009) Wootten, A. & Thompson, A. R. 2009, IEEE Proceedings, 97, 1463
- Wright et al. (2023) Wright, G. S., Rieke, G. H., Glasse, A., et al. 2023, PASP, 135, 048003
- Wright et al. (2015) Wright, G. S., Wright, D., Goodson, G. B., et al. 2015, PASP, 127, 595
- Zhang et al. (2024) Zhang, L., Packham, C., Hicks, E. K. S., et al. 2024, ApJ, 974, 195
- Zubovas & Bourne (2017) Zubovas, K. & Bourne, M. A. 2017, MNRAS, 468, 4956

### v4

@@ Acknowledgements.
- Third Reference Catalogue of Bright Galaxies. External Links: ADS entry Cited by: §1.
- SpeX Near-infrared Spectroscopic Extinction Curves in the Milky Way. ApJ 930 (1), pp. 15. External Links: Document, 2204.13716, ADS entry Cited by: §C.1.
- Modeling the Pan-Spectral Energy Distribution of Starburst Galaxies. III. Emission Line Diagnostics of Ensembles of Evolving H II Regions. ApJS 167 (2), pp. 177–200. External Links: Document, astro-ph/0608062, ADS entry Cited by: §3.5.3.
- Mid-Infrared Fine-Structure Line Ratios in Active Galactic Nuclei Observed with the Spitzer IRS: Evidence for Extinction by the Torus. ApJ 664 (1), pp. 71–87. External Links: Document, 0704.0547, ADS entry Cited by: Appendix E, §3.4.
- Observational Evidence of Active Galactic Nuclei Feedback. ARA&A 50, pp. 455–489. External Links: Document, 1204.4114, ADS entry Cited by: §1.
- Optical and mid-infrared line emission in nearby Seyfert galaxies. A&A 675, pp. A74. External Links: Document, 2301.02252, ADS entry Cited by: Figure 5, §3.3, §3.3, §3.3.
- CLOUDY 90: Numerical Simulation of Plasmas and Their Spectra. PASP 110 (749), pp. 761–778. External Links: Document, ADS entry Cited by: §3.5.2.
- Quasar feedback revealed by giant molecular outflows. A&A 518, pp. L155. External Links: Document, 1006.1655, ADS entry Cited by: §1.
- AGN wind scaling relations and the co-evolution of black holes and galaxies. A&A 601, pp. A143. External Links: Document, 1702.04507, ADS entry Cited by: §1, §1, §1.
- An Analysis of the Shapes of Interstellar Extinction Curves. VII. Milky Way Spectrophotometric Optical-through-ultraviolet Extinction and Its R-dependence. ApJ 886 (2), pp. 108. External Links: Document, 1910.08852, ADS entry Cited by: §C.1.
- Properties of the multiphase outflows in local (ultra)luminous infrared galaxies. MNRAS 505 (4), pp. 5753–5783. External Links: Document, 2006.13232, ADS entry Cited by: §1, §1.
- Cold molecular outflows in the local Universe and their feedback effect on galaxies. MNRAS 483 (4), pp. 4586–4614. External Links: Document, 1805.05352, ADS entry Cited by: §1, §1, §1.
- Star formation in southern Seyfert galaxies. MNRAS 300 (3), pp. 757–766. External Links: Document, astro-ph/9804298, ADS entry Cited by: §1.
- The KMOS3D Survey: Demographics and Properties of Galactic Outflows at z = 0.6-2.7. ApJ 875 (1), pp. 21. External Links: Document, 1807.04738, ADS entry Cited by: §1.
- The Nuclear Region of NGC 1365: Star Formation, Negative Feedback, and Outflow Structure. ApJ 913 (2), pp. 139. External Links: Document, 2104.04928, ADS entry Cited by: §1, §3.1.
- The Galaxy Activity, Torus, and Outflow Survey (GATOS). III. Revealing the inner icy structure in local active galactic nuclei. A&A 681, pp. L7. External Links: Document, 2310.09093, ADS entry Cited by: §1.
- ALMA images the many faces of the ¡ASTROBJ¿NGC 1068¡/ASTROBJ¿ torus and its surroundings. A&A 632, pp. A61. External Links: Document, 1909.00675, ADS entry Cited by: §1.
- Modeling the High-energy Ionizing Output from Simple Stellar and X-Ray Binary Populations. ApJ 960 (1), pp. 13. External Links: Document, 2307.00050, ADS entry Cited by: §3.3.
- FUSE Measurements of Far-Ultraviolet Extinction. III. The Dependence on R(V) and Discrete Feature Limits from 75 Galactic Sightlines. ApJ 705 (2), pp. 1320–1335. External Links: Document, 0909.3087, ADS entry Cited by: §C.1.
- One Relation for All Wavelengths: The Far-ultraviolet to Mid-infrared Milky Way Spectroscopic R(V)-dependent Dust Extinction Relationship. ApJ 950 (2), pp. 86. External Links: Document, 2304.01991, ADS entry Cited by: §C.1.
- Milky Way Mid-Infrared Spitzer Spectroscopic Extinction Curves: Continuum and Silicate Features. ApJ 916 (1), pp. 33. External Links: Document, 2105.05087, ADS entry Cited by: §C.1.
- dust_extinction: Interstellar Dust Extinction Models. The Journal of Open Source Software 9 (100), pp. 7023. External Links: Document, ADS entry Cited by: §C.1.
- Dusty, Radiation Pressure-Dominated Photoionization. II. Multiwavelength Emission Line Diagnostics for Narrow-Line Regions. ApJS 153 (1), pp. 75–91. External Links: Document, astro-ph/0404176, ADS entry Cited by: §3.5.3.
- Modelling the nebular emission from primeval to present-day star-forming galaxies. MNRAS 462 (2), pp. 1757–1774. External Links: Document, 1607.06086, ADS entry Cited by: §3.3.
- Probing the Excitation of Extreme Starbursts: High-Resolution Mid-Infrared Spectroscopy of Blue Compact Dwarfs. ApJ 704 (2), pp. 1159–1173. External Links: Document, 0908.4084, ADS entry Cited by: §3.3.
- The KMOS AGN Survey at High redshift (KASHz): the prevalence and drivers of ionized outflows in the host galaxies of X-ray AGN. MNRAS 456 (2), pp. 1195–1220. External Links: Document, 1511.00008, ADS entry Cited by: §1.
- Kiloparsec-scale outflows are prevalent among luminous AGN: outflows and feedback in the context of the overall AGN population. MNRAS 441 (4), pp. 3306–3347. External Links: Document, 1403.3086, ADS entry Cited by: §1.
- Observational Tests of Active Galactic Nuclei Feedback: An Overview of Approaches and Interpretation. Galaxies 12 (2), pp. 17. External Links: Document, 2404.08050, ADS entry Cited by: §1, §1.
- The European ALMA Regional Centre Network: A Geographically Distributed User Support Model. The Messenger 162, pp. 24–29. External Links: Document, 1601.04498, ADS entry Cited by: §A.3.
- MICONIC: Dual active galactic nuclei, star formation, and ionised gas outflows in NGC 6240 seen with MIRI/JWST. A&A 693, pp. A321. External Links: Document, 2412.14707, ADS entry Cited by: §1, §3.4.
- A biconical ionised gas outflow and evidence of positive feedback in NGC 7172 uncovered by MIRI/JWST. A&A 690, pp. A350. External Links: Document, 2407.15807, ADS entry Cited by: §3.4.
- Warm molecular hydrogen in outflows from ultraluminous infrared Galaxies. MNRAS 439 (3), pp. 2701–2716. External Links: Document, 1311.0311, ADS entry Cited by: §1.
- Mid-infrared Atomic Fine-structure Emission-line Spectra of Luminous Infrared Galaxies: Spitzer/IRS Spectra of the GOALS Sample. ApJ 777 (2), pp. 156. External Links: Document, 1309.4788, ADS entry Cited by: §3.3.
- The Extragalactic Distance Database: Color-Magnitude Diagrams. AJ 138 (2), pp. 332–337. External Links: Document, 0902.3675, ADS entry Cited by: §1.
- The Fornax cluster of galaxies.. MNRAS 191, pp. 685–700. External Links: Document, ADS entry Cited by: §1.
- Spatially resolved electron density in the narrow line region of z ¡ 0.02 radio AGNs. A&A 618, pp. A6. External Links: Document, 1806.02839, ADS entry Cited by: §3.4.
- SUPER. II. Spatially resolved ionised gas kinematics and scaling relations in z 2 AGN host galaxies. A&A 642, pp. A147. External Links: Document, 2008.01728, ADS entry Cited by: §1.
- Theoretical Modeling of Starburst Galaxies. ApJ 556 (1), pp. 121–140. External Links: Document, astro-ph/0106324, ADS entry Cited by: Figure 5, §3.5.3.
- The host galaxies and classification of active galactic nuclei. MNRAS 372 (3), pp. 961–976. External Links: Document, astro-ph/0605681, ADS entry Cited by: §1, Figure 5, §3.3.
- The Cosmic BPT Diagram: Confronting Theory with Observations. ApJ 774 (1), pp. L10. External Links: Document, 1307.0514, ADS entry Cited by: §3.3.
- Imaging the nuclear environment of NGC 1365 with the Hubble Space Telescope. A&A 328, pp. 483–492. External Links: ADS entry Cited by: §1.
- On the variation of the initial mass function. MNRAS 322 (2), pp. 231–246. External Links: Document, astro-ph/0009005, ADS entry Cited by: §3.5.1.
- Wavelength calibration and resolving power of the JWST MIRI Medium Resolution Spectrometer. A&A 656, pp. A57. External Links: Document, 2109.04254, ADS entry Cited by: §1, §2.
- A Note on the Viability of Gaseous Ionization in Active Galaxies by Fast Shocks. ApJ 496 (2), pp. L71–L74. External Links: Document, astro-ph/9802164, ADS entry Cited by: §3.3.
- A 3D Drizzle Algorithm for JWST and Practical Application to the MIRI Medium Resolution Spectrometer. AJ 166 (2), pp. 45. External Links: Document, 2306.05520, ADS entry Cited by: §A.1, Appendix B, §2.
- The PHANGS-JWST Treasury Survey: Star Formation, Feedback, and Dust Physics at High Angular Resolution in Nearby GalaxieS. ApJ 944 (2), pp. L17. External Links: Document, 2212.02667, ADS entry Cited by: §1.
- The PHANGS-HST Survey: Physics at High Angular Resolution in Nearby Galaxies with the Hubble Space Telescope. ApJS 258 (1), pp. 10. External Links: Document, 2101.02855, ADS entry Cited by: §1.
- Optical long slit velocities and a combined optical and H I velocity field for the barred spiral galaxy NGC 1365.. A&AS 120, pp. 403–414. External Links: ADS entry Cited by: §1.
- PHANGS-JWST First Results: Stellar-feedback-driven Excitation and Dissociation of Molecular Gas in the Starburst Ring of NGC 1365?. ApJ 944 (2), pp. L19. External Links: Document, 2212.09652, ADS entry Cited by: §A.3, §C.2, §1, §3.1, §3.1, §3.1, §3.1.
- PyNeb: a new tool for analyzing emission lines. I. Code description and validation of results. A&A 573, pp. A42. External Links: Document, 1410.6662, ADS entry Cited by: Appendix G, Figure 6, §3.4, §3.4.
- Molecular outflows in local galaxies: Method comparison and a role of intermittent AGN driving. A&A 633, pp. A134. External Links: Document, 1911.05608, ADS entry Cited by: §1, §1.
- Galaxy-scale ionised winds driven by ultra-fast outflows in two nearby quasars. A&A 644, pp. A15. External Links: Document, 2009.11294, ADS entry Cited by: §C.1.
- MIRACLE: I. Unveiling the multi-phase, multi-scale physical properties of the active galaxy NGC 424 with MIRI, MUSE, and ALMA. A&A 701, pp. A113. External Links: Document, 2503.21921, ADS entry Cited by: §A.1, Appendix B, §2, §3.6.
- MOKA3D: An innovative approach to 3D gas kinematic modelling. I. Application to AGN ionised outflows. A&A 677, pp. A58. External Links: Document, 2307.01854, ADS entry Cited by: §2, §3.6.
- Evidence of the fast acceleration of AGN-driven winds at kiloparsec scales. Nature Astronomy 9, pp. 907–915. External Links: Document, 2503.24359, ADS entry Cited by: §1, §2, §3.2, §3.6.
- HOMERUN: A new approach to photoionization modeling: I. Reproducing observed emission lines with percent accuracy and obtaining accurate physical properties of the ionized gas. A&A 689, pp. A78. External Links: Document, 2401.13028, ADS entry Cited by: §3.5.1, §3.5.2, §3.5.
- On the origin of optical and IR emission lines in star-forming galaxies. MNRAS 525 (2), pp. 2916–2934. External Links: Document, 2308.05604, ADS entry Cited by: §3.3.
- The MAGNUM survey: different gas properties in the outflowing and disc components in nearby active galaxies with MUSE. A&A 622, pp. A146. External Links: Document, 1811.07935, ADS entry Cited by: §C.2, §C.2, Appendix D, §1, §2, §3.2, §3.3, §3.4, §3.5.3.
- Exploring the Mysterious High-ionization Source Powering [Ne V] in High-z Analog SBS0335-052 E with JWST/MIRI. ApJ 985 (2), pp. 253. External Links: Document, 2502.07662, ADS entry Cited by: §3.3.
- JWST MIRI Flight Performance: Detector Effects and Data Reduction Algorithms. PASP 135 (1049), pp. 075004. External Links: Document, 2308.16327, ADS entry Cited by: §A.1.
- Star formation and AGN feedback in the local Universe: Combining LOFAR and MaNGA. A&A 665, pp. A144. External Links: Document, 2206.01195, ADS entry Cited by: §1.
- Abundance scaling in stars, nebulae and galaxies. MNRAS 466 (4), pp. 4403–4422. External Links: Document, 1612.03546, ADS entry Cited by: §3.5.2.
- Astrophysics of gaseous nebulae and active galactic nuclei. External Links: ADS entry Cited by: §C.2, Appendix D, Appendix G, §3.4.
- Geometric distortion and astrometric calibration of the JWST MIRI Medium Resolution Spectrometer. A&A 682, pp. A53. External Links: Document, 2307.01025, ADS entry Cited by: §A.1.
- No evidence of active galactic nucleus features in the nuclei of Arp 220 from JWST/NIRSpec IFS. A&A 690, pp. A171. External Links: Document, 2403.13948, ADS entry Cited by: §1, §3.6.
- Updated point spread function simulations for JWST with WebbPSF. In Space Telescopes and Instrumentation 2014: Optical, Infrared, and Millimeter Wave, J. M. Oschmann, M. Clampin, G. G. Fazio, and H. A. MacEwen (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9143, pp. 91433X. External Links: Document, ADS entry Cited by: Appendix B.
- The diverse cold molecular gas contents, morphologies, and kinematics of type-2 quasars as seen by ALMA. A&A 658, pp. A155. External Links: Document, 2111.13578, ADS entry Cited by: §1.
- JWST MIRI reveals the diversity of nuclear mid-infrared spectra of nearby type 2 quasars. A&A 698, pp. A194. External Links: Document, 2504.01595, ADS entry Cited by: §3.4, §3.4.
- Optical and JWST Mid-IR Emission Line Diagnostics for Simultaneous IMBH and Stellar Excitation in z 0 Dwarf Galaxies. ApJ 927 (2), pp. 165. External Links: Document, 2202.01330, ADS entry Cited by: §3.3.
- Radiative cooling of swept-up gas in AGN-driven galactic winds and its implications for molecular outflows. MNRAS 478 (3), pp. 3100–3119. External Links: Document, 1710.09433, ADS entry Cited by: §1.
- The origin of fast molecular outflows in quasars: molecule formation in AGN-driven galactic winds. MNRAS 474 (3), pp. 3673–3699. External Links: Document, 1706.03784, ADS entry Cited by: §1.
- The Mid-Infrared Instrument for the James Webb Space Telescope, I: Introduction. PASP 127 (953), pp. 584. External Links: Document, 1508.02294, ADS entry Cited by: §1.
- Active galactic nuclei winds as the origin of the H2 emission excess in nearby galaxies. MNRAS 491 (1), pp. 1518–1529. External Links: Document, 1909.11742, ADS entry Cited by: §1.
- Highly Ionized Iron Absorption Lines from Outflowing Gas in the X-Ray Spectrum of NGC 1365. ApJ 630 (2), pp. L129–L132. External Links: Document, astro-ph/0508608, ADS entry Cited by: §1.
- Detection of CO Hot Spots Associated with Young Clusters in the Southern Starburst Galaxy NGC 1365. ApJ 654 (2), pp. 782–798. External Links: Document, astro-ph/0609805, ADS entry Cited by: §1.
- Measuring mass: Non-circular motions of gas in disk galaxies and radial velocities of stars in a globular cluster. Ph.D. Thesis, Rutgers University, New Jersey. External Links: ADS entry Cited by: §1.
- The central region of NGC 1365. VLA and SEST observations of the radio continuum and CO.. A&A 295, pp. 585. External Links: ADS entry Cited by: §1.
- PHANGS-JWST First Results: Rapid Evolution of Star Formation in the Central Molecular Gas Ring of NGC 1365. ApJ 944 (2), pp. L15. External Links: Document, 2212.09168, ADS entry Cited by: §C.2, §1.
- Three-Dimensional Integral Field Observations of 10 Galactic Winds. I. Extended Phase (gsim10 Myr) of Mass/Energy Injection Before the Wind Blows. ApJ 711 (2), pp. 818–852. External Links: Document, 1001.4315, ADS entry Cited by: §1.
- Multiphase characterization of AGN winds in five local type-2 quasars. A&A 681, pp. A63. External Links: Document, 2311.10132, ADS entry Cited by: §1.
- Re-evaluating old stellar populations. MNRAS 479 (1), pp. 75–93. External Links: Document, 1805.08784, ADS entry Cited by: §3.5.1.
- The complex radio and X-ray structure in the nuclear regions of the active galaxy NGC 1365. MNRAS 306 (2), pp. 479–490. External Links: Document, astro-ph/9902206, ADS entry Cited by: §1.
- Kinematics and dynamics of the central region in the barred galaxy NGC 1365.. MNRAS 221, pp. 1. External Links: Document, ADS entry Cited by: §1.
- Connecting X-ray nuclear winds with galaxy-scale ionised outflows in two z 1.5 lensed quasars. A&A 648, pp. A99. External Links: Document, 2102.07789, ADS entry Cited by: §C.1.
- JWST/NIRSpec insights into the circumnuclear region of Arp 220: A detailed kinematic study. A&A 693, pp. A36. External Links: Document, 2407.08505, ADS entry Cited by: §1, §3.6.
- UV-extended E-MILES stellar population models: young components in massive early-type galaxies. MNRAS 463 (4), pp. 3409–3436. External Links: Document, 1612.01187, ADS entry Cited by: §C.2.
- Galactic Winds. ARA&A 43 (1), pp. 769–826. External Links: Document, astro-ph/0504435, ADS entry Cited by: §1.
- Spectral Classification of Emission-Line Galaxies. ApJS 63, pp. 295. External Links: Document, ADS entry Cited by: §1.
- MAGNUM survey: Compact jets causing large turmoil in galaxies. Enhanced line widths perpendicular to radio jets as tracers of jet-ISM interaction. A&A 648, pp. A17. External Links: Document, 2011.04677, ADS entry Cited by: §1.
- Complex AGN feedback in the Teacup galaxy. A powerful ionised galactic outflow, jet-ISM interaction, and evidence for AGN-triggered star formation in a giant bubble. A&A 678, pp. A127. External Links: Document, 2309.02498, ADS entry Cited by: §1.
- Ionized gas outflows from the MAGNUM survey: NGC 1365 and NGC 4945. Frontiers in Astronomy and Space Sciences 4, pp. 46. External Links: Document, 1801.05448, ADS entry Cited by: §1, §2, §3.2.
- MAGNUM survey: A MUSE-Chandra resolved view on ionized outflows and photoionization in the Seyfert galaxy NGC1365. A&A 619, pp. A74. External Links: Document, 1809.01206, ADS entry Cited by: §C.2, §C.2, Appendix D, §1, §1, §1, §1, §2, §3.1, §3.1, §3.2, §3.3, §3.3, §3.4, §3.6, 1st item.
- A catalogue of quasars and active nuclei: 12th edition. A&A 455 (2), pp. 773–777. External Links: Document, ADS entry Cited by: §1.
- Mid-infrared Properties of the Swift Burst Alert Telescope Active Galactic Nuclei Sample of the Local Universe. I. Emission-line Diagnostics. ApJ 716 (2), pp. 1151–1165. External Links: Document, 1004.5321, ADS entry Cited by: §3.3.
- The Prevalence of Gas Outflows in Type 2 AGNs. ApJ 817 (2), pp. 108. External Links: Document, 1511.05142, ADS entry Cited by: §1.
- The Atacama Large Millimeter/Submillimeter Array. IEEE Proceedings 97 (8), pp. 1463–1471. External Links: Document, 0904.3739, ADS entry Cited by: §1.
- The Mid-Infrared Instrument for the James Webb Space Telescope, II: Design and Build. PASP 127 (953), pp. 595. External Links: Document, 1508.02333, ADS entry Cited by: §1.
- The Mid-infrared Instrument for JWST and Its In-flight Performance. PASP 135 (1046), pp. 048003. External Links: Document, ADS entry Cited by: §1, §1.
- The Galaxy Activity, Torus, and Outflow Survey (GATOS). IV. Exploring Ionized Gas Outflows in Central Kiloparsec Regions of GATOS Seyferts. ApJ 974 (2), pp. 195. External Links: Document, 2409.09771, ADS entry Cited by: §1, §3.4.
- Do AGN outflows quench or enhance star formation?. MNRAS 468 (4), pp. 4956–4967. External Links: Document, 1703.10782, ADS entry Cited by: §1.